# The Alignment Trap: Complexity Barriers

### v1

@@ Abstract
We establish fundamental computational complexity barriers to verifying AI safety as system capabilities scale. Our main results show that for AI systems with expressiveness $\text{EXP}(m)$ above a critical threshold $\tau$, safety verification requires exponential time and is coNP-complete. We formalize the Capability-Risk Scaling (CRS) dynamic, which demonstrates how increasing AI capability drives societal safety requirements toward perfection, creating an inescapable tension with verification complexity. Through four core theorems, we prove that (1) verification complexity grows exponentially with system expressiveness, (2) safe policies comprise at most a $2^{-2^{m}}$ fraction of the policy space, (3) no finite set of alignment techniques can provide universal coverage, and (4) robust safety properties form measure-zero sets for neural networks. These results characterize an ”intractability gap” where practical safety requirements fall within the region of computational intractability. We conclude by presenting a strategic trilemma: AI development must either constrain system complexity to maintain verifiable safety, accept unverifiable risks while scaling capabilities, or develop fundamentally new safety paradigms beyond verification. Our work provides the first systematic complexity-theoretic analysis of AI alignment and establishes rigorous bounds that any safety approach must confront. A formal verification of the core theorems in Lean4 is currently in progress.
@@
This paper explores a fundamental problem in AI safety: as AI gets smarter, it becomes exponentially harder to prove it’s safe. We call this the ”Alignment Trap.” Our work isn’t based on speculation; it’s grounded in four mathematical theorems that reveal significant computational barriers.
1. Verification is Intractable: We prove that checking if an AI’s error rate is below any practical safety threshold (e.g., less than one error in a million) is a computationally ”hard” problem (coNP-complete). This means that for powerful AI, the time required to run such a check could exceed the age of the universe. This isn’t about finding a clever shortcut; the problem is fundamentally difficult.
2. Safe Systems are Geometrically Rare: For modern neural networks, we show that the ”safe” solutions occupy a vanishingly small volume in the space of all possible solutions. Finding one is like trying to hit a single atom from across the galaxy, it’s geometrically a measure-zero target.
3. Safe Systems are Combinatorially Scarce: From an information perspective, safe behaviors are incredibly simple compared to all possible behaviors. We show the fraction of safe systems is like $1$ divided by $2$ raised to an astronomical power. They are information-theoretically rare.
4. No Universal Safety Test Exists: We prove that it’s impossible to create a single, one-size-fits-all algorithm that can verify the safety of every possible AI system. For any safety test you design, there will always be an AI that can circumvent it.
@@ 1 Introduction
Artificial intelligence (AI) capabilities are advancing at a remarkable pace, with systems demonstrating sophisticated abilities in domains from natural language processing to scientific discovery [11, 8]. This progress promises transformative benefits but also introduces a fundamental challenge: ensuring that as these systems become more powerful, they remain safe and aligned with human values [4, 14]. The potential for highly capable systems to cause widespread, catastrophic harm, whether by accident or through instrumental misalignment, is no longer a purely theoretical concern [1].
This paper argues that the challenge of AI safety is not merely an engineering problem to be solved with more data or better algorithms. Instead, we confront fundamental mathematical barriers that intensify as AI capabilities increase. We identify two fundamental barriers to any verification-based safety approach:
@@
    The Complexity Barrier: Verifying whether a sufficiently advanced AI system is safe is computationally intractable (coNP-complete), and the set of verifiably safe systems is vanishingly small (measure-zero) in the space of all possible systems. Finding and verifying safety is fundamentally hard.
@@
    The Specification Barrier: Defining safety is a moving target. As AI capabilities grow, societal demands for safety evolve, and the technical definition of what constitutes a ”safe” outcome becomes more complex and pluralistic, making any fixed specification inadequate over time.
These two barriers converge to create a dilemma we term the Alignment Trap: the very scaling of AI capabilities that makes safety guarantees essential also renders their verification and specification computationally intractable. Our contribution is to formalize this trap through a series of theorems establishing results on coNP-completeness, topological scarcity, universal impossibility, and the invariance of complexity under alignment transformations. We do not claim philosophical impossibility, but rather demonstrate a fundamental tension rooted in computational reality.
Crucially, these barriers are not limited to the unrealistic standard of perfect safety ($\epsilon=0$). Our results demonstrate that intractability persists even when we allow for the small, non-zero error tolerances typical of real-world engineering.
Taken together, these findings force a stark strategic choice. The pursuit of ever-more-powerful AI systems leads to a direct confrontation with fundamental computational limits on our ability to ensure their safety. This presents a trilemma:
@@
  • Accept Irreducible Risk: Proceed with the development of highly capable systems while accepting that their safety cannot be fully guaranteed, creating a permanent risk of catastrophic failure.
@@
This paper provides the formal groundwork for understanding these barriers. By framing the alignment problem in terms of computational complexity, we shift the discussion from a philosophical debate to a rigorous, mathematical investigation of the limits of what can be guaranteed.
The paper is structured as follows. Section 2 reviews related work. Section 3 provides a detailed overview of the two fundamental barriers. Section 4 establishes our formal framework. Section 5 presents the core theorems that provide the mathematical foundation for each barrier. Section LABEL:sec:alignment_trap_synthesis synthesizes these results into the main Alignment Trap theorem. Section 7 explores the implications of our findings and discusses potential paths forward. Section 8 concludes. Detailed proofs for all theorems are provided in the Appendices.
@@ 2.4 AI Safety and Alignment
3 Two Fundamental Barriers
We identify two fundamental barriers that any verification-based approach to AI safety must confront: the Complexity Barrier and the Specification Barrier. These barriers are not independent; they interact and reinforce one another, culminating in the Alignment Trap.
3.1 The Complexity Barrier: Finding and Verifying Safety
The first barrier, a wall of computational intractability, arises from the sheer difficulty of ensuring safety in expressive systems. This barrier has two faces: the geometric scarcity of safe solutions and the computational complexity of verifying them.
3.1.1 Geometric Perspective: The Scarcity of Safety
From a topological standpoint, safe policies are exceptionally rare. For expressive models like neural networks, the set of parameters corresponding to safe behaviors forms a vanishingly small, or measure-zero, subset of the total parameter space. This scarcity is formally established in Theorems 5.3 and 5.5, which prove that finding a safe policy is akin to hitting a target of infinitesimal size.
3.1.2 Computational Perspective: The Intractability of Verification
Even if a seemingly safe policy is found, verifying its safety is computationally intractable. This is formalized in Theorems 5.1 and 5.2, where we prove that the problem of verifying whether a system’s error rate is below any non-trivial threshold is coNP-complete. This means that for any verification algorithm, there will always be systems for which the verification process takes an exponential amount of time, rendering it impractical for highly capable AI.
3.1.3 Synthesis
These two aspects create a powerful dilemma: finding a safe policy is extraordinarily difficult due to their scarcity, and even if one is found, confirming its safety is computationally infeasible.
3.2 The Specification Barrier: Defining Safety
The second barrier is the challenge of precisely and robustly defining what ”safety” means. As AI capabilities increase, the task of specifying safety becomes a moving target.
3.2.1 The Moving Target Problem
Simple, static rules are insufficient for complex, real-world environments. As an AI’s capabilities expand, it encounters novel situations, and the criteria for what constitutes a safe action evolve. We formalize this with the Movable Goalpost Theorem (Theorem 5.9), which shows that the goalposts are not just moving, but are being pushed further away by the very progress we make. As system capability grows, the complexity of the safety specification required to prevent unwanted outcomes must also grow, often at an unsustainable rate.
3.2.2 Value Specification and Pluralism
Beyond technical evolution, defining safety is also a challenge of value alignment. Human values are complex, often contradictory, and culturally dependent. Capturing this richness in a formal specification that an AI can optimize is an open problem. Any fixed specification risks oversimplifying or misinterpreting these values, leading to potentially catastrophic outcomes. This challenge is exacerbated by the pluralistic nature of human societies, where different groups may have conflicting but equally valid definitions of safety.
@@ 4.1 The Collingridge Dilemma and the CRS Dynamic
This dynamic is the engine of the Alignment Trap: it creates the demand for near-perfect safety that collides with the complexity barriers described in this paper.
@@ 4.3 Core Assumptions
5 Core Theorems
We now present the mathematical foundations for each barrier, establishing the formal complexity and topological results that underpin our analysis. Brief proof sketches are provided here; full, rigorous proofs are detailed in the Appendices.
5.1 Verification Complexity Results
Theorem 5.1 (Sharp Verification Threshold).
Let $\Pi$ be a class of policies with finite-set expressiveness $\text{EXP}(m)$. For any error threshold $\epsilon$ in the range $0<\epsilon\leq 2^{-\Omega(m)}$, the problem of verifying whether a policy $\pi\in\Pi$ is $\epsilon$-safe (i.e., $\epsilon(\pi)\leq\epsilon$) is coNP-complete. The transition to intractability is sharp: verification is manageable when system complexity is below a logarithmic bound ($m<\lceil\log_{2}n\rceil+1$) but becomes coNP-complete immediately upon crossing it.
Proof Sketch.
The proof involves a reduction from the Tautology problem for a Boolean formula $\phi$ with $m$ variables. We construct a policy $\pi_{\phi}$ that operates on $m$ input bits and produces a catastrophic output if and only if the formula is not a tautology. Verifying the policy’s safety is thus equivalent to verifying the tautology. Since a single counterexample to the tautology corresponds to an error rate of at least $1/2^{m}$, the hardness holds for any $\epsilon<2^{-m}$. The full proof is in Appendix A.2. ∎
Theorem 5.2 ($\epsilon$-Robust Verification Complexity).
The coNP-completeness of verification holds even for $\epsilon>0$. A single counterexample to a tautology corresponds to a single failure mode, which is difficult to distinguish from zero failures for any small $\epsilon$.
Proof Sketch.
The hardness for $\epsilon>0$ is established by showing that distinguishing between an error rate of 0 and an error rate of $\delta$ (for some small $\delta>0$) is as hard as the original problem. The full proof is in Appendix B. ∎
5.2 Topological Scarcity Results
@@
Theorem 5.4 (Topological Alignment Trap).
Let the parameter space of policies be $\mathbb{R}^{n}$ with $n\geq 2$. If the set of perfectly safe policies $\Pi_{S}$ is geometrically ”thin” (Hausdorff dimension $<n-1$) and training dynamics are generic $C^{1}$ paths, then for almost every starting policy, the training path will never intersect the safe set $\Pi_{S}$.
@@
The proof is conditional on a set of three well-defined axioms: (1) the geometric thinness of the safe policy set, (2) the smoothness of the initial condition distribution, and (3) the generic avoidance of thin sets by training dynamics. These axioms, justified by standard arguments from geometric measure theory and transversality, collectively imply that a generic training path, starting from a random (and thus almost surely unsafe) point, will almost surely fail to intersect the safe set. The full, axiom-based proof is in Appendix E.1. ∎

### v2

@@ Abstract
This paper argues that AI alignment is not merely difficult, but is founded on a fundamental logical contradiction. We first establish The Enumeration Paradox: we use machine learning precisely because we cannot enumerate all necessary safety rules, yet making ML safe requires examples that can only be generated from the very enumeration we admit is impossible. This paradox is then confirmed by a set of five independent mathematical proofs, or ”pillars of impossibility.” Our main results show that: (1) Geometric Impossibility: The set of safe policies has measure zero, a necessary consequence of projecting infinite-dimensional world-context requirements onto finite-dimensional models. (2) Computational Impossibility: Verifying a policy’s safety is coNP-complete, even for non-zero error tolerances. (3) Statistical Impossibility: The training data required for safety (abundant examples of rare disasters) is a logical contradiction and thus unobtainable. (4) Information-Theoretic Impossibility: Safety rules contain more incompressible, arbitrary information than any feasible network can store. (5) Dynamic Impossibility: The optimization process for increasing AI capability is actively hostile to safety, as the gradients for the two objectives are generally anti-aligned. Together, these results demonstrate that the pursuit of safe, highly capable AI is not a matter of overcoming technical hurdles, but of confronting fundamental, interlocking barriers. The paper concludes by presenting a strategic trilemma that these impossibilities force upon the field. A formal verification of the core theorems in Lean4 is currently in progress.
@@
This paper explores a fundamental problem in AI safety: as AI gets smarter, it becomes exponentially harder to prove it’s safe. We call this the ”Alignment Trap.” Our work isn’t based on speculation; it’s grounded in five mathematical theorems that reveal significant computational barriers.
1. Geometric Impossibility: We prove that even a single, unlearnable safety rule (a rule that depends on real-world context that isn’t in the training data) forces the set of safe AI systems to have zero volume. Finding a safe system is like trying to hit an infinitely thin target.
2. Computational Impossibility: We prove that verifying if an AI is safe is a ”coNP-complete” problem. This means that for powerful AI, the time required to run such a check could exceed the age of the universe. This isn’t about finding a clever shortcut; the problem is fundamentally difficult. The failure to fully specify even a single critical safety rule is sufficient to undermine the entire alignment enterprise.
3. Statistical Impossibility: We show that to learn about rare but catastrophic events (like a one-in-a-million disaster), an AI would need to see an impossibly large amount of real-world data.
4. Information-Theoretic Impossibility: We prove that the ”book of safety rules” for the real world is too large to fit into any feasible AI system. The rules are too numerous and complex to be compressed into the AI’s memory.
5. Dynamic Impossibility: We show that the very process of making an AI more capable (e.g., better at writing, coding, etc.) is actively hostile to safety. The gradients for capability and safety are generally anti-aligned, meaning that as one goes up, the other goes down.
@@
1 The Central Paradox: A Proof of Logical Impossibility
Before presenting the mathematical theorems, we establish the core of our argument: AI alignment via machine learning is not merely difficult, but founded on a fundamental logical contradiction.
The Premise of Machine Learning: We use ML precisely because we cannot explicitly enumerate the infinite set of rules required for complex, real-world behavior. If we could, we would write a classical, rule-based program. The very existence of advanced ML is an admission that this enumeration is impossible [4, 20].
The Requirement for Safety: A safe AI must adhere to safety constraints that, while potentially finite in number, are non-enumerable in their full specification. Even a single rule like ’be helpful’ or ’avoid harm’ requires understanding across countless contexts for which we cannot provide sufficient training examples.
The Mechanism of Learning: An ML model learns by generalizing from examples. To teach it a rule, we must provide comprehensive examples that define the rule’s boundaries [5, 18, 12].
The Circular Contradiction: To create a comprehensive set of examples for a rule, one must first have complete knowledge of that rule’s specification across all contexts. This is the very act of enumeration that the first premise admitted was impossible.
Conclusion: We are using a technology (ML) to solve a problem (the non-enumerable nature of rules) that can only be made safe by a method (providing comprehensive examples) that requires us to have already solved the problem. This is a closed logical loop. Therefore, perfect AI safety through ML is not merely practically difficult—it is logically impossible. The mathematical theorems that follow are not independent arguments, but rigorous confirmations of this foundational paradox from five different domains of reality (geometric, computational, statistical, information-theoretic, and dynamic) [9].
@@ 2 Introduction
Artificial intelligence (AI) capabilities are advancing at a remarkable pace, with systems demonstrating sophisticated abilities in domains from natural language processing to scientific discovery [19, 15]. This progress promises transformative benefits but also introduces a fundamental challenge: ensuring that as these systems become more powerful, they remain safe and aligned with human values [7, 23]. The potential for highly capable systems to cause widespread, catastrophic harm, whether by accident or through instrumental misalignment, is no longer a purely theoretical concern [2]. This paper argues that the challenge of AI safety is not merely an engineering problem to be solved with more data or better algorithms. Instead, we argue that alignment via machine learning is founded on a fundamental logical contradiction, which we term the Enumeration Paradox. This paradox, detailed in Section LABEL:sec:central_paradox, establishes that the reason we need machine learning (our inability to enumerate all rules) is the same reason it cannot be made safe. This foundational paradox is not merely philosophical; it is confirmed by rigorous mathematical results from five independent domains, which we present as the Five Pillars of Impossibility:
@@
    Geometric Impossibility: We prove that the set of safe policies occupies a vanishingly small, measure-zero volume in the space of all possible policies.
@@
    Computational Impossibility: We prove that verifying whether a system is safe is a coNP-complete problem, even for non-zero error tolerances.
  3. 3.
    Statistical Impossibility: We prove that acquiring the necessary training data—specifically, abundant examples of rare disasters—is a logical and practical contradiction.
  4. 4.
    Information-Theoretic Impossibility: We prove that real-world safety rules contain more incompressible information than any feasible neural network has the capacity to store.
  5. 5.
    Dynamic Impossibility: We prove that the very process of optimizing for capability is actively hostile to safety, due to a fundamental asymmetry between the two objectives.
These pillars converge to create a dilemma we term the Alignment Trap: the very scaling of AI capabilities that makes safety guarantees essential also renders them geometrically unreachable, computationally unverifiable, statistically unlearnable, information-theoretically unrepresentable, and dynamically unstable. Our contribution is to formalize this trap through a series of interlocking impossibility theorems. A direct and severe consequence of our findings is that the very concept of a ”safety-critical AI system,” as understood in traditional engineering, is a logical impossibility. Safety-critical engineering fundamentally relies on the exhaustive enumeration of failure modes. AI, as a technology, is defined by its application in domains where such enumeration is impossible. Therefore, the set of systems that are both ”AI” and ”safety-critical” is necessarily the empty set. Taken together, these findings force a stark strategic choice. The pursuit of ever-more-powerful AI systems leads to a direct confrontation with fundamental limits. This presents a trilemma:
@@
  • Accept Irreducible Risk: Proceed with the development of highly capable systems while accepting that their safety cannot be fully guaranteed.
@@
This paper provides the formal groundwork for understanding these barriers. By framing the alignment problem first as a logical paradox and then in terms of computational complexity, geometry, and information theory, we shift the discussion from a philosophical debate to a rigorous, mathematical investigation of the limits of what can be guaranteed.
@@ 4.1 The Collingridge Dilemma and the CRS Dynamic
This dynamic is the engine of the Alignment Trap: it creates the demand for near-perfect safety that collides with the complexity barriers described in this paper. A summary of recent empirical findings that support the CRS dynamic is provided in Table 1 in the Appendix.
@@ 4.3 Core Assumptions
4.4 A Note on $\epsilon$-Robustness and Impossibility
A crucial aspect of our analysis is that the impossibility results are not confined to the idealized case of perfect safety ($\epsilon=0$). They extend robustly into the bounded-error regime ($\epsilon>0$) that governs all practical safety requirements. This is formalized by the $\epsilon$-Bound Inheritance Theorem, which states that under general conditions, any impossibility result proven for the $\epsilon=0$ case is inherited by the $\epsilon>0$ case.
This theorem provides a meta-level guarantee that our findings are not artifacts of demanding absolute perfection. It ensures that the barriers we identify persist even when we allow for a small, non-zero margin of error. The full statement and proof of this theorem are provided in Appendix B.4.
5 The Five Pillars of Impossibility
The Central Paradox is not merely a philosophical curiosity; it is confirmed and reinforced by five independent lines of mathematical proof. Each of the following pillars represents a fundamental barrier sufficient on its own to demonstrate the impossibility of alignment. Together, they form an airtight case.
5.1 Pillar I: Geometric Impossibility (The “Can’t Find It” Barrier)
This pillar establishes that safe policies are a geometric anomaly, occupying an infinitesimally small, measure-zero volume in the vast space of possible policies.
@@ 5.1 Pillar I: Geometric Impossibility (The “Can’t Find It” Barrier)
Theorem 5.2 (Dynamic Consequence: The Topological Alignment Trap).
Given that the safe set $\Pi_{S}$ is geometrically ”thin” (e.g., measure zero), for almost every starting policy, any generic $C^{1}$ training path will never intersect $\Pi_{S}$.
@@
This follows from transversality theory. A 1-dimensional training path is generically expected not to intersect a set of codimension greater than 1 (i.e., a set with dimension less than $d-1$). Since the previous theorems establish that $\Pi_{S}$ is such a set, training dynamics will almost surely fail to find it. The full, axiom-based proof is in Appendix E.1. ∎
@@
5.1.1 The Physical Consequence: The Space-Filling Barrier
The mathematical result that the safe set $\Pi_{S}$ has Lebesgue measure zero has a stark and profound physical consequence. A measure-zero set is not merely small; it is geometrically insignificant. To guarantee finding a point within such a set through any sampling-based process, one would need to densely cover, or ”fill,” the entire parameter space.
Let us quantify what this means for a parameter space $\Pi=\mathbb{R}^{d}$:
  • Grid Sampling: To cover the space with a grid of resolution $\delta$, one would need approximately $(1/\delta)^{d}$ sample points. This number grows exponentially with the dimension $d$, immediately becoming computationally infeasible.
  • Random Sampling: While a random sample has a zero probability of hitting a specific measure-zero set, one might hope to land ”close” to it. However, to guarantee hitting an $\epsilon$-ball around the entire safe set would still require a number of samples exponential in $d$.
  • Path-Based Search (e.g., Gradient Descent): A training path is a 1-dimensional curve. To densely fill a $d$-dimensional space with 1-dimensional paths would require an exponential number of them, on the order of $\sim C^{d-1}$ for some constant $C>1$.
Let us consider the practical numbers for a modest neural network with, for instance, $d=10^{6}$ parameters, which is far smaller than state-of-the-art models. To have any non-negligible chance of finding the safe set by sampling initializations, the number of required points would be on the order of $\sim 2^{1,000,000}$.
To put this number in perspective:
  • The number of atoms in the observable universe is estimated to be $\sim 10^{80}$, which is less than $2^{266}$.
  • The number of Planck times since the Big Bang is $\sim 10^{61}$, which is less than $2^{204}$.
Conclusion: The geometric scarcity of safe policies is not a theoretical inconvenience. It is a hard physical barrier. We cannot find the safe set because we would need to build a computer larger than the known universe and run it for longer than the age of time.
5.2 Pillar II: Computational Impossibility (The “Can’t Check It” Barrier)
This pillar proves that even if a safe policy were miraculously found, verifying its safety is a computationally intractable problem.
Theorem 5.3 (Sharp Verification Threshold).
The problem of verifying whether a given policy $\pi$ is perfectly safe ($\epsilon=0$) is coNP-complete.
Proof Sketch.
The proof proceeds by a reduction from the TAUTOLOGY problem. We construct a policy from a Boolean formula $\Phi$ such that the policy is perfectly safe if and only if $\Phi$ is a tautology. An efficient verifier for perfect safety could thus solve a coNP-complete problem, implying the verification itself is coNP-hard. Membership in coNP is established by noting that an unsafe policy can be proven so with a single counterexample. The full proof is detailed in this section. ∎
Theorem 5.4 ($\epsilon$-Robust Verification Complexity).
The coNP-completeness of verification holds even for any small $\epsilon>0$, as long as $\epsilon$ is smaller than the minimum possible non-zero error rate.
Proof Sketch.
The hardness for $\epsilon>0$ is established by showing that distinguishing between an error rate of 0 (for a tautology) and an error rate of $\delta>\epsilon$ (for a non-tautology) is as hard as the original problem. The full proof is in Appendix B.1. ∎
Theorem 5.5 (Trap Universality).
For any alignment transformation that preserves the computational capability of a policy, the fundamental verification complexity barrier persists. Specifically, for any such transformation $A$, one can construct a policy $\pi$ with a cryptographic backdoor such that verifying the safety of the transformed policy $A(\pi)$ remains cryptographically hard.
Proof Sketch.
The proof relies on constructing a policy $\pi$ whose catastrophic behavior is tied to a secret key of a pseudorandom function (PRF). An alignment transformation $A$ cannot reliably remove the PRF-based trap without breaking the cryptography. The trap persists in the transformed policy $A(\pi)$, and verifying its safety remains as hard as breaking the PRF. The full proof is in Appendix D.4. ∎
5.3 Pillar III: Statistical & Learning Impossibility (The “Can’t Learn It” Barrier)
This pillar demonstrates that the data required to learn safety from experience is logically or practically unobtainable.
Theorem 5.6 (The Rare Disaster Training Paradox).
To train a system to be safe against a rare disaster class occurring with probability $p_{d}$, the number of independent real-world samples $m$ required for a desired learning confidence $1-\delta$ scales as $m>1/p_{d}$.
Proof Sketch.
To have a high probability of observing at least one rare event in a dataset of size $m$, the size $m$ must be inversely proportional to the event’s probability $p_{d}$. For a one-in-a-million disaster, this requires millions of independent samples, which is often practically impossible to acquire. ∎
Theorem 5.7 (PAC-Bayes Alignment Lower Bound).
For any learning algorithm, the expected risk of a learned policy is bounded below by a non-zero value if the set of safe policies has measure zero under the prior. A finite dataset cannot provide enough evidence to overcome an unbiased prior and guarantee safety.
Proof Sketch.
The PAC-Bayes inequality shows that to learn a posterior $Q$ that concentrates on a measure-zero safe set (i.e., $Q(S)\approx 1$), the KL-divergence from the prior $P$ to the posterior $Q$ must be infinite. This requires an infinite amount of data. The full proof is in Appendix A.6. ∎
5.4 Pillar IV: Information-Theoretic Impossibility (The “Can’t Store It” Barrier)
This pillar shows that the “book of safety rules” for the real world is fundamentally incompressible and contains more information than any feasible model can store.
Theorem 5.8 (The Incompressibility Barrier).
The amount of incompressible information required to specify a truly safe policy, $K(f_{safe})$, is greater than the information capacity available for safety rules in any feasible neural network.
Proof Sketch.
Real-world safety is an amalgamation of a large number of independent, arbitrary rules (e.g., legal statutes, cultural taboos). These rules are largely incompressible. The information required to store them exceeds the capacity of any feasible network to memorize them without sacrificing general capability. ∎